OpenStudy (anonymous):

[2.06] Solve the equation Quantity x minus 5 all over 3 = 6 for x. 13 14 23 10

terenzreignz (terenzreignz):

\[\Large \frac{x-5}3 = 6\]

terenzreignz (terenzreignz):

First thing you want to do is get rid of that denominator...

OpenStudy (anonymous):

ya

terenzreignz (terenzreignz):

So... how do you go about that? You multiply both sides of the equation by 3. Now what happens when you do that?

OpenStudy (anonymous):

x-15=18

terenzreignz (terenzreignz):

Almost... except 5 should not have been multiplied by 3. The only thing multiplying 3 does to the left-side is get rid of the denominator. Try again...?

OpenStudy (anonymous):

x-5=18

terenzreignz (terenzreignz):

Right... so now, you get rid of that -5 in the left-side...

OpenStudy (anonymous):

23

terenzreignz (terenzreignz):

Good job :)